A chemist needs to strengthen a 15% alcoholic solution to one of 32% alcohol .how much pure alcohl should be added to 800 ml of 15% solution?
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

A chemist needs to strengthen a 15% alcoholic solution to one of 32% alcohol .how much pure alcohl should be added to 800 ml of 15% solution?
---------
Add L liters
800*15 + L*100 = (800 + L)*32

Answer by greenestamps(13200)



Here is an alternative to the standard algebraic solution method that requires only a few simple calculations.

(1) Find how far the desired 32% is from the 15% of the original solution and from the 100% of the pure alcohol.
100-32 = 68; 32-15 = 17

(2) The ratio of the two numbers from (1) is the ratio in which the two ingredients must be mixed.
68:17 = 4:1

Since the 32% is closer to 15% than to 100%, you need 4 parts of the original solution to 1 part of the added pure alcohol. In other words, the amount of pure alcohol to be added should be 1/4 the amount of the original mixture.
1/4 of 800ml = 200ml

Answer: He needs to add 200ml of pure alcohol.